OpenStudy (anonymous):

4) Simplify 4 / √5 − 1 by rationalizing the denominator

OpenStudy (anonymous):

multiply numerator and denominator by the "conjugate" of the denominator \[\frac{4}{\sqrt{5}-1}\times \frac{\sqrt{5}+1}{\sqrt{5}+1}\]

OpenStudy (anonymous):

denominator is \[5-1=4\] because \[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

4 root5 + 1 4(root5 + 1) 4(root5 +1) -------1 x --------- = ---------------- = ------------ = root5 + 1 root5 -1 root5 + 1 (root5)^2 - (1)^2 5 - 1

OpenStudy (anonymous):

then in this case you can cancel the 4 but not usually \[\frac{4(\sqrt{5}+1)}{4}=\sqrt{5+1}\]

OpenStudy (anonymous):

final ans is \[\sqrt{5} + 1\] and NOT \[\sqrt{5+1}\]

OpenStudy (anonymous):

@satellite I think by mistake u put the 1 also under the root sign.....